# Upsample or Upweight? Balanced Training on Heavily Imbalanced Datasets

### v3

@@ 1 Introduction
We closely investigate this assumed equivalency in theory (§3.1). Specifically, we prove that although they are equivalent in full gradient descent (Theorem 1), Temperature Sampling induces lower variance in the context of stochastic gradient descent (Theorem 2). Moreover, the variance induced by scalarization increases as the approximated temperature increases or the domain distribution’s skewness increases (Theorem 3). Based on our theoretical results and connecting to the literature on lower variance between stochastic gradients accelerates convergence Sutskever et al. (2013); mccandlish2018empiricalmodellargebatchtraining, we make the following hypothesis:
@@ 5 Related Works
The search space for the optimal domain weights at any given training iteration is combinatorically large. Existing works conduct comprehensive experiments on smaller scaled models to learn how the training and generalization error varies according to dataset sizes and domain weights —— “scaling laws" of domain mixture Ye et al. (2024); ge2024bimixbivariatedatamixing; jiang2024adaptivedataoptimizationdynamic. Closer to our work, Chen et al. (2023) fits scaling laws for sampling temperature for multilingual machine translation. Concurrent to our work, he2024scalinglawsmultilinguallanguage fits scaling laws for multilingual language modeling. However, as jiang2024adaptivedataoptimizationdynamic denotes: “the optimal data policy for a smaller model does not necessarily generalize to larger models Albalak et al. (2023); Ye et al. (2024)."
@@ Acknowledgements
This work is supported by a generous gift the Allen Institute for AI and We are grateful Nicholas Lourie and Jingyu Zhang for their insightful feedback throughout this project. GPU machines for conducting experiments were provided by the DSAI cluster.

### v4

@@ 1 Introduction
We closely investigate this assumed equivalency in theory (§3.1). Specifically, we prove that although they are equivalent in full gradient descent (Theorem 1), Temperature Sampling induces lower variance in the context of stochastic gradient descent (Theorem 2). Moreover, the variance induced by scalarization increases as the approximated temperature increases or the domain distribution’s skewness increases (Theorem 3). Based on our theoretical results and connecting to the literature on lower variance between stochastic gradients accelerates convergence Sutskever et al. (2013); McCandlish et al. (2018), we make the following hypothesis:
@@ 5 Related Works
The search space for the optimal domain weights at any given training iteration is combinatorically large. Existing works conduct comprehensive experiments on smaller scaled models to learn how the training and generalization error varies according to dataset sizes and domain weights — “scaling laws" of domain mixture Ye et al. (2024); Ge et al. (2024); Jiang et al. (2024). Closer to our work, Chen et al. (2023) fits scaling laws for sampling temperature for multilingual machine translation. Concurrent to our work, He et al. (2024) fits scaling laws for multilingual language modeling. However, as Jiang et al. (2024) denotes: “the optimal data policy for a smaller model does not necessarily generalize to larger models Albalak et al. (2023); Ye et al. (2024)."
@@ Acknowledgements
This work is supported by ONR grant (N00014-24-1-2089) and a gift from Allen Institute for AI. We are grateful Nicholas Lourie and Jingyu Zhang for their insightful feedback throughout this project. GPU machines for conducting experiments were provided by the DSAI cluster.